An electrical firm manufactures light bulb that has a length of life that is approximately normally distributed with a mean of 800 hours and a standard deviation of 40 hours. The supervising electrical Engineer took a random sample of 30 bulbs with an average life of 788 hours, test the hypothesis that µ=800 hours against the alternative hypothesis µ is greater than 800.



Use confidence level of 96%.



Is it a two tailed or one tailed test?



Null hypothesis?



Alternative hypothesis?



Size of the test?



Test statistics?



Critical region?



Decision ( accept or reject null hypothesis)?

1
Expert's answer
2022-05-25T15:56:09-0400

The following null and alternative hypotheses need to be tested:

"H_0:\mu=800"

"H_1:\mu>800"

This corresponds to a one-tailed right-tailed test, for which a z-test for one mean, with known population standard deviation will be used.

Size of the test is "n=30."

Based on the information provided, the significance level is "\alpha = 0.04," and the critical value for a right-tailed test is "z_c = 1.7507."

The rejection region for this right-tailed test is "R = \{z:z>1.7507\}."

The z-statistic is computed as follows:


"z=\dfrac{\bar{x}-\mu}{\sigma/\sqrt{n}}=\dfrac{788-800}{40/\sqrt{30}}\approx-1.6432"

Since it is observed that "z=-1.6432<1.7507=z_c," it is then concluded that the null hypothesis is not rejected.

Using the P-value approach:

The p-value is "p=P(z>-1.6432)=0.949829," and since "p=0.949829>0.04=\alpha," it is concluded that the null hypothesis is not rejected.

Therefore, there is not enough evidence to claim that the population mean "\mu"

is greater than 800, at the "\alpha = 0.04" significance level.